OpenStudy (anonymous):

Can you help me find the local extrema and intervals of concavity for g(x)=(x^2+2x-15)/(x^2+2x-3)?

OpenStudy (anonymous):

Re-write the derivative please. I forgot ahaha ! :D

OpenStudy (anonymous):

g'(x)=24(x+1)/(x+3)^2(x-1)^2 .... do I begin by plugging in all critical points into the original formula to get the local max and min?

OpenStudy (anonymous):

For Extrema, only use the critical numbers where derivative is Zero. that is x = -1

OpenStudy (anonymous):

but that's only one number?

OpenStudy (anonymous):

I need max AND min

OpenStudy (anonymous):

So what, you can't always have both :D As for y=x^2 You have only one minimum and NO Maximum

OpenStudy (anonymous):

Now you have to check where x=-1 is minimum or maximum

OpenStudy (anonymous):

and how do I check? like I understand how to get max and min and apparently there is no max in this one... so what i'm asking is how to I check to make sure that i'm right ?

OpenStudy (anonymous):

You have to check for second derivative :D if its positive you have minimum if its negative you have max

OpenStudy (anonymous):

I have to plug -1 into the original equation right? and not into the derivative

OpenStudy (anonymous):

What do you mean by take the 2nd derivative>? as in why do I do that

OpenStudy (anonymous):

In this case, it is ambiguous that x = -1 is min or max. Right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So to remove this ambiguity, second derivative comes to the rescue :D

OpenStudy (anonymous):

so I plugged -1 into the original equation of g(x)=x^2+2x-15/x^2+2x-3 and got -4

OpenStudy (anonymous):

Yeah the point of extremum is (-1,-4) but you dont know its maxx or min.. Find second derivative to confirm that !

OpenStudy (anonymous):

Ok, be right back

OpenStudy (anonymous):

g''(x)=-24[3x+6x+7]/(x+3)^3(x-1)^3 is the second derivative

OpenStudy (anonymous):

[3x^2+6x+7]

OpenStudy (anonymous):

Ok, Plug in -1 :) and tell it's positive or negative

OpenStudy (anonymous):

alright, be right back again then haha

OpenStudy (anonymous):

undefined?

OpenStudy (anonymous):

Hey No !! :/ It is positive.. indicating local minimum and -1

OpenStudy (anonymous):

So I may have just calculated it wrong then, eh?

OpenStudy (anonymous):

Unfortunately :P Why is it undefined.. :D check again

OpenStudy (anonymous):

division by 0

OpenStudy (anonymous):

It's -1 not 1 :D There is no Zero in denominator

OpenStudy (anonymous):

yeah i've been putting -1 and it keeps coming with the denominator as 0

OpenStudy (anonymous):

\[g''(-1) = \frac{-24(3(-1)^2+6(-1)+7)}{(-1+3)^3 (-1-1)^3}\] Just calculate this now

OpenStudy (anonymous):

96?

OpenStudy (anonymous):

Whatever be the magnitude.. just check the sign.. Here it is positive right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so this would be the max? and the -4 would be the min?

OpenStudy (anonymous):

Now how you got 96 ? Why it isnt still undefined ? :D

OpenStudy (anonymous):

Since g''(x) is positive.. x=-1 is a MINIMUM for g(x) Hence there is only one minimum at (-1,-4) and no Maximum for g(x)

OpenStudy (anonymous):

alright, makes sense

OpenStudy (anonymous):

-4 is the Y-Value of g(x) at x = -1 !! Got it ? Cheers !

OpenStudy (anonymous):

yes, thank you!

OpenStudy (anonymous):

i have to go now Bye See You Later :D Fan me if you want help from me again ;)

OpenStudy (anonymous):

Bye. Thank you again!!!!!